# From Prototypes to General Distributions: An Efficient Curriculum for Masked Image Modeling

### v1

@@ 4.2 Main Results
The results demonstrate the clear advantages of a prototype-driven curriculum. Most notably, our approach achieves 47.40% nearest neighbor classification accuracy with 800 epochs of training, substantially outperforming the 800-epoch (30.25%) MAE baseline. Interestingly, while [35] found that non-prototypical (hard) samples need to be preserved when reducing dataset size in supervised learning, our results reveal this strategy to be suboptimal for self-supervised MIM training. The significant performance gap between hard-sample curriculum (62.09% LP, 82.95% FT) and our prototypical approach (68.84% LP, 83.31% FT) suggests that early exposure to prototypical examples plays a crucial role in developing robust visual representations. This aligns with our hypothesis that building from prototypical to complex examples creates a more stable optimization trajectory in the challenging MIM setting.
@@
Figure 5: Training curve. Neareast-neighbor accuracy throughout training for an MAE model trained with and without a prototype-driven curriculum.
@@
4.3 Impact of Temperature Annealing
@@
Table 3: Impact of temperature on training. Performance comparison between fixed temperature sampling and our annealed approach. The temperature parameter $\tau$ controls the sampling distribution, with lower values favoring prototypical examples and $\tau=\infty$ corresponding to uniform sampling over the entire dataset. $|{\mathcal{D}}_{\tau}|/|{\mathcal{D}}|$ is the effective dataset size. In the annealed experiment, the temperature $\tau$ is adjusted from 0.05 to 0.2.
Temperature annealing enables the model to transition from prototypical to diverse examples during training. To examine its role in our curriculum strategy, we compared it against variants using fixed temperatures $\tau$ ranging from 0.01 to infinity in Tab. 3. Constant temperature effectively creates static training distributions, with lower temperatures prioritizing prototypical examples, and higher temperatures emphasizing diverse samples.
@@
(b) DiNO backbone architecture.
# Clusters ${\mathcal{D}}{\mathcal{B}}$ Optimal NN LP
@@
(c) Number of clusters.
Table 4: Empirical study on prototype identification. We evaluate the impact of (a) pre-training methods, (b) model architectures, and (c) number of clusters. We further compare self-supervised DiNO clusters against “Oracle” clusters which use ground truth labels to define the centroids. Unless otherwise specified, experiments use DiNO features from a ViT-Base/16 backbone with 1000 clusters. All models are trained for 800 epochs on ImageNet-1K.

### v2

@@ 4.2 Main Results
The results demonstrate the clear advantages of a prototype-driven curriculum. Most notably, our approach achieves 47.40% nearest neighbor classification accuracy with 800 epochs of training, substantially outperforming the 800-epoch (30.25%) MAE baseline.
Interestingly, while [40] found that non-prototypical (hard) samples need to be preserved when reducing dataset size in supervised learning, our results reveal this strategy to be suboptimal for self-supervised MIM training. The significant performance gap between hard-sample curriculum (62.09% LP, 82.95% FT) and our prototypical approach (68.84% LP, 83.31% FT) suggests that early exposure to prototypical examples plays a crucial role in developing robust visual representations. This aligns with our hypothesis that building from prototypical to complex examples creates a more stable optimization trajectory in the challenging MIM setting.
@@
Figure 5: Training curve. Nearest-neighbor accuracy throughout training for an MAE model trained with and without a prototype-driven curriculum.
@@
Few-shot Performance
To further validate the quality and effectiveness of the representations learned by our accelerated pretraining method, we perform linear probing evaluations under a few-shot scenario. Specifically, we assess the performance of our pretrained models on eight fine-grained datasets, using a 5-shot setup where only five labeled examples per class are available for training the linear classifier. This evaluation tests the generalization capabilities and data efficiency of the learned representations. Results are summarized in Tab. 3, which clearly demonstrate that our method, even at significantly fewer training epochs (200 and 400 epochs), achieves comparable or superior performance compared to the baseline trained for 1600 epochs, highlighting both efficiency and effectiveness with up to an 8$\times$ speedup.
CUB-200 [45] SUN397 [48] OxfordPets [33] Caltech101 [13] UCF101 [39] DTD [8] Flowers102 [30] Food101 [4] AVG
Baseline (1600 epoch) 14.42 ±0.38 30.20 ±0.61 39.76 ±2.43 74.81 ±2.11 49.06 ±0.67 44.47 ±1.89 64.48 ±0.85 16.97 ±0.42 41.77
Ours (200 epoch) 14.35 ±0.41 30.15 ±1.03 43.79 ±1.75 77.42 ±1.51 48.15 ±1.05 42.81 ±1.79 64.13 ±0.71 16.64 ±0.05 42.18
Ours (400 epoch) 18.46 ±0.40 34.33 ±1.24 56.97 ±2.20 82.02 ±1.86 52.35 ±1.47 45.59 ±1.20 68.99 ±0.53 19.11 ±0.36 47.23
Table 3: 5-shot linear probing accuracy (%) on eight fine-grained datasets. Mean and standard deviation reported over three runs. Our accelerated method (200 and 400 epochs) achieves comparable or improved performance compared to the baseline method (1600 epochs), highlighting a training speedup of up to 8×.
@@
Table 4: Impact of temperature on training. Comparison between fixed and annealed temperature sampling. Lower temperatures ($\tau$) favor prototypical examples; $\tau=\infty$ represents uniform sampling. $|{\mathcal{D}}_{\tau}|/|{\mathcal{D}}|$ denotes effective dataset size. Annealed sampling adjusts $\tau$ from 0.07 to 0.6.
@@
(b) DiNO Backbone.
# Clusters NN LP
@@
(c) Number of clusters. 978 is computed via DB 3.2.1.
Table 5: Empirical study on prototype identification. We evaluate the impact of (a) feature space, (b) model architectures, and (c) number of clusters. We further compare self-supervised DiNO clusters against “Oracle” clusters which use ground truth labels to define the centroids. Unless otherwise specified, experiments use DiNO features from a ViT-Base/16 backbone with 1000 clusters. All models are trained for 800 epochs on ImageNet-1K.
4.3 Impact of Temperature Annealing
Temperature annealing enables the model to transition from prototypical to diverse examples during training. To examine its role in our curriculum strategy, we compared it against fixed temperature $\tau$ schedules, with temperatures ranging from 0.01 to infinity in Tab. 5. Constant temperature effectively creates static training distributions, with lower temperatures prioritizing more prototypical examples.
@@ 5 Conclusion
Supplementary Material
The supplementary material provides a detailed configuration used for downstream tasks, as well as additional visualizations of the prototypical scores obtained across different feature spaces.
Training Configuration for Downstream Tasks.
[…]
Visualization of various Feature Spaces.
We visualize the images with various distances from their respective cluster centroid in different feature spaces. The visualization is shown in Figure 6. The DINO representations space (regardless of the size of the backbone architecture) consistently places prototypical views of each object closer to the cluster centroids. Results using low-level features, such as MAE or SIFT, appear to be less organized.
Parameter Value
Optimizer AdamW
Training epochs 100
Batch size 1024
Base learning rate (LR) 1e-3
LR schedule cosine decay
LR Warm-up epochs 5
Weight decay 0.05
Optimizer momentum $\beta_{1},\beta_{2}{=}0.9,0.999$
Layer-wise LR decay 0.75
Augmentation RandAug (9, 0.5)
Label smoothing 0.1
Mixup 0.8
CutMix 1.0
Drop path 0.1
Table 6: End-to-end fine-tuning setting.
Parameter Value
Optimizer LARS
Training epochs 90
Batch size 16384
Base learning rate (LR) 0.1
LR schedule cosine decay
LR Warm-up epochs 10
Weight decay 0
Optimizer momentum 0.9
Augmentation RandomResizedCrop
Table 7: Linear probing setting.
(a) Visualization for synsets n01677366 (iguana).
(b) Visualization for synsets n01829413 (hornbill).
(c) Visualization for synsets n02097658 (silky terrier).
Figure 6: Visualization of cluster samples arranged by their distances from their respective cluster centers (Part 1 of 3). Each row represents images from the same model, while each column indicates the distance of images from their cluster’s centroid.
(a) Visualization for synsets n02268853 (damselfly).
(b) Visualization for synsets n02490219 (marmoset).
(c) Visualization for synsets n02747177 (trash can).
Figure 7: Visualization of cluster samples arranged by their distances from their respective cluster centers (Part 2 of 3). Each row represents images from the same model, while each column indicates the distance of images from their cluster’s centroid.
(a) Visualization for synsets n03110669 (cornet).
(b) Visualization for synsets n03785016 (moped).
(c) Visualization for synsets n03796401 (moving van).
Figure 8: Visualization of cluster samples arranged by their distances from their respective cluster centers (Part 3 of 3). Each row represents images from the same model, while each column indicates the distance of images from their cluster’s centroid.